Malita has K150 of disposable income to spend each week and cannot borrow money. She buys Milk balls and a composite good. Suppose that Milk balls cost K2.50 per bag and the composite good costs K1 per unit.Sketch Malita’s budget constraint.What is the opportunity cost, in terms of bags of Milk balls, of an additional unit of the composite good?Suppose that in an inflationary period the cost of the composite good increases to K1.50 per unit, but the cost of Milk balls remains the same. Sketch the new budget constraint.What is the opportunity cost of van additional unit of the composite good?Suppose now Malita demands a pay raise to figjht the inflation. Her boss submits and raises her salary so that her disposable income is now K225 per week. Sketch the new budget constraint. Is Malita better off?What is the opportunity cost of an additional unit of the composite good?


1
Expert's answer
2021-06-16T13:36:30-0400

1)

The equation of budget line will be:

(Price of milk balls ×quantity of milk balls) + (price of composite good ×Quantity of composite good)=Income

"2.50M+1C=150"

where M refers to quantity purchased of milk balls and C refers to quantity purchased of composite good. 

Now, we will find the intercept of both the axis using the below formula. 

Intercept of X-axis÷maximum quantity consumed of milk balls=income÷price of one milk ball

"=\frac{150}{2.50}\\=60"

Intercept of y-axis÷maximum quantity consumed of composite good=income÷price of one unit of composite good

"=\frac{150}{1}\\=150"

Now, we will take quantity of milk balls on x-axis and quantity of composite goods on the y-axis to plot the budget constraint. We have represented the budget constraint by line AB



To calculate the opportunity cost of composite good, we will divide the price of composite good by the price of milk balls. 

Opportunity cost of composite good=price of composite good÷price of milk balls

"\frac{1}{2.5}\\=0.4"

This implies that to consume one more unit of composite good, 0.4 milk ball bags need to be sacrificed. 


2)If the cost of the composite good increases to K1.50 per unit, the new budget constraint will be:

The equation of budget line will be:

(Price of milk balls ×quantity of milk balls) + (price of composite good ×Quantity of composite good)=Income

"2.50M+1.5C=150"

The intercept will also change.

Intercept of X-axis÷maximum quantity consumed of milk balls=income÷price of one milk ball

"=\frac{150}{2.50}\\=60"

Intercept of y-axis÷maximum quantity consumed of composite good=income÷price of one unit of composite good

"=\frac{150}{1.5}\\=100"

The price of milk balls remains the same and only the price of composite good increases. So, the budget line will pivot. The x-axis intercept will remain the same and the y-axis intercept will fall. The new budget line is represented by BC



To calculate the opportunity cost of composite good, we will divide the price of composite good by the price of milk balls. 

Opportunity cost of composite good=price of composite good÷price of milk balls

"\frac{1.5}{2.5}\\=0.6"

This implies that to consume one more unit of composite good, 0.6 milk ball bags need to be sacrificed.


3)

The equation of budget line will be:

(Price of milk balls ×quantity of milk balls) + (price of composite good ×Quantity of composite good)=Income

"2.50M+1.5C=225"

The intercept will also change.

Intercept of X-axis÷maximum quantity consumed of milk balls=income÷price of one milk ball

"=\frac{225}{2.50}\\=90"

Intercept of y-axis÷maximum quantity consumed of composite good=income÷price of one unit of composite good

"=\frac{225}{1.5}\\=150"

The new budget line is represented by AC




Malita better off.

To calculate the opportunity cost of composite good, we will divide the price of composite good by the price of milk balls. 

Opportunity cost of composite good=price of composite good÷price of milk balls

"\frac{1.5}{2.5}\\=0.6"

This implies that to consume one more unit of composite good, 0.6 milk ball bags need to be sacrificed.